OpenStudy (anonymous):

What is 2xy/y^4 and (-x(y^4))/y^4?

OpenStudy (anonymous):

\[2xy/y^4\] and \[-xy^4/y^4\] to be clear.

OpenStudy (anonymous):

what do u want to do with them? solve them simultaneously or differentiate them implicitly

OpenStudy (anonymous):

?

OpenStudy (anonymous):

generally they are implicit functions... and you can do many things with them even derive a differential equation...

OpenStudy (anonymous):

Sorry, for the late reply. I just want to divide them. I'm pretty bad when it comes to this. What do you get when you divide 2xy by y^4 and similarly for the second one?

OpenStudy (anonymous):

\[2xy/y^4 \div (-x(y^4))/y^4\]

OpenStudy (anonymous):

\[2xy /y^4= 2x/y^3\]

OpenStudy (anonymous):

Ok, so would the first one be \[2xy/y^4 = 2xy^-3\] ?

OpenStudy (anonymous):

\[\frac{2xy}{y^4}= \frac{2x}{y^3}\]

OpenStudy (anonymous):

what happened is

OpenStudy (anonymous):

Cool, same thing then. I'm pretty sure I get it. Thanks!

OpenStudy (anonymous):

when u divide u subtract the power

OpenStudy (anonymous):

when u multiply u add the powers

OpenStudy (anonymous):

Got it. I just wasn't certain if the x was affected in any way. Thanks again.